The height of a triangle is twice the base. The area is 9 square inches. find the area.
Answer by fractalier(6550)

The area is given as 9 -- is that what they really want? I doubt it...
The facts they give us are
h = 2b
and
A = 9
but we know that
A = (1/2)bh
so now substitute in everything and get
9 = (1/2)b(2b)
b^2 = 9
b = 3
Then h = 6.
This is probably what they wanted.